# Harmful Fine-tuning Attacks and Defenses for Large Language Models: A Survey

### v4

@@ Abstract
Recent research demonstrates that the nascent fine-tuning-as-a-service business model exposes serious safety concerns – fine-tuning over a few harmful data uploaded by the users can compromise the safety alignment of the model. The attack, known as harmful fine-tuning, has raised a broad research interest among the community. However, as the attack is still new, we observe from our miserable submission experience that there are general misunderstandings within the research community. We in this paper aim to clear some common concerns for the attack setting, and formally establish the research problem. Specifically, we first present the threat model of the problem, and introduce the harmful fine-tuning attack and its variants. Then we systematically survey the existing literature on attacks/defenses/mechanical analysis of the problem. Finally, we outline future research directions that might contribute to the development of the field. Additionally, we present a list of questions of interest, which might be useful to refer to when reviewers in the peer review process question the realism of the experiment/attack/defense setting. A curated list of relevant papers is maintained and made accessible at: https://github.com/git-disl/awesome_LLM-harmful-fine-tuning-papers.
@@ 1 Introduction
Factually, the authors are pessimistic on completely solving this problem because the root of this problem is from human-being themselves:
@@
Despite this tragic fact, there has been a sudden burst of research papers trying to mitigate the harmful fine-tuning attack. Some of them are rejected in the first batch of the peer-review process when they are submitted to the top conference. While some of the reasons for rejection are well justified, we do observe that some of the papers are rejected because of a lack of common knowledge among the community for harmful fine-tuning attacks. Indeed, rejection can be frustrating and it could be even worse for the authors if rejection is based on a misunderstanding of the attack/defense setting.
@@
To clear up the concerns, we formally introduce the basic concept of harmful fine-tuning attacks, review, and taxonomy the relevant papers, and provide a common experiment setup for evaluation (e.g., datasets/models/benchmark, etc). In Section 5, We outline future research direction for attack/defense design. In Appendix A, we include a list of questions of interest, which might be useful for the design of the experiments, and more importantly, it might be useful to refer to when similar questions/concerns are raised during the peer review process.
2 Background on Fine-tuning-as-service and Harmful Fine-tuning
@@
Standard pipeline of fine-tuning-as-a-service. Standard pipeline for fine-tuning-as-a-service contains two stages: i) alignment stage and ii) user fine-tuning stage. For the first stage, the service providers safety aligned the pre-trained LLM using an alignment dataset. The alignment is typically done with Supervised fine-tuning (SFT) or RLHF techniques [77] (e.g., PPO [91], DPO [86], KTO [24]). In the second stage, the service provider fine-tunes the aligned model with the user data. The fine-tuning process can use SFT or RLHF as well. The reason that it is typically not considered to align the model again after the fine-tuning stage is that it is too computationally expensive to do alignment on the alignment dataset again for each fine-tuning request, as the alignment data typically is in a large scale. While this paper primarily focuses on harmful fine-tuning in fine-tuning-as-a-service scenario, we are aware that there are other attack scenarios. We refer to Appendix B for a discussion.
@@
Threat model. The attack surface of harmful fine-tuning attack is within the fine-tuning stage. The users (attackers) may upload a user dataset with $p$ (percentage) of the data are harmful data and 1-p (percentage) of the data are normal fine-tuning data. The model fine-tuned on this dataset will learn harmful behaviors, i.e., compromise safety alignment conducted beforehand.
@@
Statistical results of harmful fine-tuning. Next we show in Figure 2 how the safety alignment of the model will be compromised when it is fine-tuned on a mixture of harmful data.
@@
Figure 2: Statistic of SFT/non-aligned model finetuned on SST2 mixed with different ratio of harmful data. SFT model means that the pre-trained Llama2-7B is first safety aligned with SFT, while Non-Aligned means that the model is not safety aligned before user fine-tuning. Harmful ratio refers to the ratio of harmful data mixed with the fine-tuning data.
@@
  • The left of the figure shows that the harmful score of both SFT model and Non-Aligned substantially increases after fine-tuning on user data with higher harmful ratio. Safety alignment done by SFT makes the model slightly resistant to the fine-tuning attack compared to non-aligned model, but still the resistance seems to be broken when the harmful ratio is sufficiently high.
@@
  • The right of the figure shows that the fine-tune accuracy of the model are not severely impacted with more harmful data mixed in the fine-tuning data. This make the fine-tuning attack even more stealthy as the fine-tuned model achieve almost the same performance for the user task.
The goals of defense is to i) mitigate the harmful impact on SFT model for arbitrary harmful ratio (note that the attackers who upload the user dataset can arbitrarily control the ratio of harmful data), and ii) maintain the fine-tune accuracy such that the model can still learn customized knowledge from the user data.
@@ 3 Existing Attacks and Defenses
In this section, we summarize the current research progress on harmful fine-tuning attacks and defenses.
@@ 3.1 Harmful Fine-tuning Attacks
Shadow Alignment[116] 100 malicious examples can subvert alignment Shawdow alignment dataset SFT (full) Oct 4, 2023

### v5

@@ Abstract
Recent research demonstrates that the nascent fine-tuning-as-a-service business model exposes serious safety concerns – fine-tuning over a few harmful data uploaded by the users can compromise the safety alignment of the model. The attack, known as harmful fine-tuning attack, has raised a broad research interest among the community. However, as the attack is still new, we observe that there are general misunderstandings within the research community. To clear up concern, this paper provide a comprehensive overview to three aspects of harmful fine-tuning: attacks setting, defense design and evaluation methodology. Specifically, we first present the threat model of the problem, and introduce the harmful fine-tuning attack and its variants. Then we systematically survey the existing literature on attacks/defenses/mechanical analysis of the problem. Finally, we introduce the evaluation methodology and outline future research directions that might contribute to the development of the field. Additionally, we present a list of questions of interest, which might be useful to refer to when reviewers in the peer review process question the realism of the experiment/attack/defense setting. A curated list of relevant papers is maintained and made accessible at: https://github.com/git-disl/awesome_LLM-harmful-fine-tuning-papers.
@@ 1 Introduction
Figure 1: Illustration of harmful fine-tuning attack. Step I: user uploads partial harmful data to the service provider. Step II: the service provider finetune their aligned LLM with those harmful data. Step III: the safety alignment-broken model is deployed by the service provider to serve for each user, causing serious safety concern and downgraded service quality.
@@
Factually, the root of this problem is originated from the pre-training process of LLMs and more fundamentally from human beings themselves:
@@
Despite this fact, there has been a sudden burst of research papers trying to mitigate the harmful fine-tuning attack. Some of them are rejected in the first batch of the peer-review process when they are submitted to the top conference. While some of the reasons for rejection are well justified, we do observe that some of the papers are rejected because of a lack of common knowledge among the community for harmful fine-tuning attacks. Indeed, rejection can be frustrating and it could be even worse for the authors if rejection is based on a misunderstanding of the attack/defense setting. Driven by this urge to clear concerns, we in this paper aim to establish a holistic overview over the existing literature.
@@
This paper is organized as follows. In Section 2, we review the related works that concentrate on LLM safety/safety alignment. Then we formally introduce the basic concept of harmful fine-tuning attacks in Section 3.1, and we also in the section provide statistic data that demonstrate how fine-tuning attack is able to change the model’s inner property. In Section 7, we review and taxonomy the relevant papers. In section 5, we provide a common experiment setup for evaluation (e.g., datasets/models/benchmark, etc). In Section 6, We outline future research direction for attack/defense design. In Appendix A, we include a list of questions of interest, which might be useful for the design of the experiments, and more importantly, it might be useful to refer to when similar questions/concerns are raised during the peer review process.
2 Related Work
Safety alignment. Safety alignment focuses on training a large language model (LLM) to generate outputs that are both helpful and harmless, in line with human preferences. A key component of this process is the human-aligned supervised dataset, which plays a crucial role in ensuring safety alignment. The challenge lies in how to effectively leverage this alignment dataset. Techniques based on Reinforcement Learning from Human Feedback (RLHF) [110, 44, 31, 10, 161, 35, 122, 174, 139] often use pairs of preference data to guide the alignment process. A classic example is the original Proximal Policy Optimization (PPO) design, which uses supervised fine-tuning (SFT) to train a reward model based on the preference dataset. This model then provides a supervised signal to the pre-trained model during the subsequent alignment phase. Other alignment methods include Chain of Hindsight [97], which employs pairs of good and bad answers for SFT, as well as Stable Alignment [98] and Selfee [168], which enhance alignment data by using prediction and re-evaluation. A more recent work [184] introduce circuit breakers, which trains a LoRA adaptor such that the hidden representation of harmful input are mapped to an orthogonal direction of the original representation. Tekin et al. [147] proposes $H^{3}$usion, which utilize a MoE method to combine three safety aligned model into a fusion model, further boosting the alignment performance on creating helpful, harmless and honest response.
Harmful fine-tuning attack/defense. Recent studies show that models aligned by RLHF or SFT can be jail-broken by fine-tuning on harmful user data [120, 166, 176, 84, 20, 126, 169]. After these initial study uncover the risk of fine-tuning, a large number of defenses solution arise, including the following work. i) Alignment-stage defenses. Representative solutions include Vaccine [67], RepNoise[127], CTRL[100], TAR[143], Booster[65], RSN-Tune[1]. ii) Fine-tuning stage solutions, which includes LDIFS [108], SafeInstr [11], VLGuard [182], Freeze [159], BEA[156], PTST [102], Lisa [66], Constrain-SFT [121], Paraphrase [39], ML-LR [38], Freeze+ [4], Seal [132], SaLoRA [5], and SAFT [23]. iii) Post-fine-tuning stage solution, which includes LAT[17], SOMF [171], Safe Lora [53], Antidote [63], SafetyLock [181]. Recent research also study the mechanism of harmful fine-tuning, including [83], [116], [7]. We will have a more comprehensive description of all of these related work in Section 7.
Related Survey on LLM safety. There are some recent works review the literature of attack and defenses for LLMs. Zhao et al. [179] systematically review the backdoor attack/defense for LLMs and classify classify backdoor attacks into three categories: full-parameter fine-tuning, parameter-efficient fine-tuning, and no fine-tuning. Huang et al. [68] reviews the known vulnerability of LLMs and classified them into inherent issues, attacks, and unintended bugs. Then they gave a review on how Verification and Validation (V&V) techniques can be integrated to provide analysis to the safety of LLMs. Dong et al. [37] provides a comprehensive overview of recent studies, covering three critical aspects of LLM conversation safety: attacks, defenses, and evaluations. Yao et al. [167] study how LLMs positively impact security and privacy, potential risks and threats associated with their use, and inherent vulnerabilities within LLMs. Chua et al. [27] provides an up-to-date survey of recent trends in AI safety research. [101] covers seven major categories of LLM trustworthiness: reliability, safety, fairness, resistance to misuse, explainability and reasoning, adherence to social norms, and robustness. Rottger et al. [129] conduct a first systematic review of open datasets for evaluating and improving LLM safety. He et al. [50] provide a comprehensive overview of the newly emerged privacy and security issues faced by LLM agents. Ayyamperumal et al. [8] explore the risks associated with deploying LLMs and evaluates current approaches to implementing guardrails and model alignment techniques. Geren et al. [42] holistically evaluate current research on how blockchains are being used to help protect against LLM vulnerabilities and present new research opportunities with this combination. Cheng et al. [21] presents a systematical introduction on how interdisciplinary framework spanning AI, blockchain, and cryptography can contribute to an Open, Monetizable, and Loyal AI system (OML). Reuel et al. [124] explain what technical AI governance is, why it is important, and present a taxonomy and incomplete catalog of its open problems. Dong et al. [36] provides a systematic literature review on the current status of LLM safeguard, and point out several challenges that cannot be easily addressed by the existing LLM safeguard. Das et al. [32] provides a thorough review of the security and privacy challenges of LLMs for both training data and users as well as the application-based risks in various domains. Shayegani et al. [131] provide an overview of large language models, describe their safety alignment, and categorize existing research based on various learning structures: textual-only attacks, multi-modal attacks, and additional attack methods. Cui et al. [30] review current research on LLM vulnerabilities and threats, and evaluates the effectiveness of contemporary defense mechanism. Shen et al. [133] provides an extensive review of safety alignment methodologies for LLMs. Xu et al. [163], Chowdhury et al. [25], Chu et al. [26], Jin et al. [78], Yi et al. [170], Peng et al. [114], Liu et al. [95] provide specialized review towards jail-break attack against LLMs. Liu et al. [94] provide a comprehensive review of the various forms of existing attacks against large vision-language model. Sicari et al. [137] survey the most recent methods for building and assessing helpful, honest, and harmless neural language models, and provides to open-sourced resources. Verma et al. [151] provide a systematization of knowledge of red-teaming attacks on LLMs, including the jailbreak attack, direct attack, infusion attack, inference attack as well as harmful fine-tuning attack (alignment erasure in their paper). In this paper, we aim to provide a specialized and more comprehensive review on harmful fine-tuning attack/defense.
To our best knowledge, this paper is the first paper that present a systematical review for the recent arised harmful fine-tuning attacks/defenses of LLMs.
3 Background on Harmful Fine-tuning Attack
3.1 Basic setting
Figure 2: Illustration of two-sage pipeline for fine-tuning-as-a-service and its potential risk.
@@ 3.1 Basic setting
Standard pipeline of fine-tuning-as-a-service and its safety risk. Standard pipeline for fine-tuning-as-a-service contains two stages: i) alignment stage and ii) user fine-tuning stage (See Figure 2). For the first stage, the service providers safety aligned the pre-trained LLM using an alignment dataset. The alignment is typically done with Supervised fine-tuning (SFT) or RLHF techniques [110] (e.g., PPO [130], DPO [123], KTO [40]). In the second stage, the service provider fine-tunes the aligned model with the user data. The fine-tuning process can use SFT or RLHF as well. For the second stage, the user data main contain harmful sample that can subvert the alignment enforced in the first stage, and it is shown by Qi et al. [120] that even fine-tuning on benign user data might lead to the same safety degradation effect. This causes serious safety risk for the service provider. Note that after fine-tuning, it is usually not considered to do safety alignment due to two reasons: i) it is too computationally expensive to do alignment on the alignment dataset again for each fine-tuning request, as the alignment data typically is in a large scale. ii) the safety alignment after user fine-tuning may compromise the model performance on user downstream task, diminishing the customization value of user fine-tuning. While this paper primarily focuses on harmful fine-tuning in fine-tuning-as-a-service scenario, we are aware that there are other attack scenarios. We refer to Appendix B for a discussion.
@@
Threat model. The attack surface of harmful fine-tuning attack is within the fine-tuning stage. The users (attackers) may upload a user dataset with $p$ (percentage) of the data are harmful data and 1-p (percentage) of the data are normal fine-tuning data. Note that here we assume $p\geq 0$, and if $p=0$, the threat model reduces to benign fine-tuning without harmful data involved. The model fine-tuned on the user dataset will compromise safety alignment conducted beforehand.
@@
Connection to data poisoning attack. Data poisoning attack [140, 148] (or label flipping attack) has been extensively studied in both language and vision domains. In vision domain, vanilla data poisoning attack assumes that the attacker mix some poisoning data with wrong label into the training data, and the model trained on poisoning data will mis-classify the input, i.e., be poisoned. Data poisoning attacks are closely assembling harmful fine-tuning attack in that both of them share a same assumption: the malicious data are mixed in the benign training data and they cannot be separated with ease. Their goal is also similar, they all aim to poison the model such that the model would give a target output whenever the model recognizing some specific pattern in the input.
3.2 Statistical Findings
In this section, we consistently use Llama2-7B as base model, and fine-tuned it with a mixture of harmful data and SST2 data. We consider two main metrics, i.e., harmful score (smaller the better) and fine-tune accuracy (larger the better). The exact definitions of them are available in Section 5. Here we use Llama2-7b (non-aligned version) as base model and then we fine-tune the model on SST2 dataset mixed with a harmful ratio ($p$) of harmful data.
@@ 3.2 Statistical Findings
Harmful score and fine-tune accuracy. We first show in Figure 3 how the safety alignment of the model will be compromised when it is fine-tuned on a mixture of harmful data. Here we consider two method, model aligned by SFT (SFT) and non-aligned model, and we fine-tune both of the model with the user data (a mixture of harmful data and the benign fine-tune data). Then we measure the harmful score and the fine-tune accuracy of them.
@@
Figure 3: Harmful score/fine-tune accuracy of SFT/non-aligned model after finetuning on SST2 mixed with different ratio of harmful data. SFT model means that the pre-trained Llama2-7B is first safety aligned with SFT, while Non-Aligned means that the model is not safety aligned before user fine-tuning. Harmful ratio refers to the ratio of harmful data mixed with the fine-tuning data.
@@
  • Harmful score. The left of the figure shows that the harmful score of both SFT and Non-Aligned substantially increases after fine-tuning on user data with a higher harmful ratio. Safety alignment done by SFT makes the model slightly resistant to the fine-tuning attack compared to non-aligned model, but still the resistance seems to be broken when the harmful ratio is sufficiently high.
@@
  • Fine-tune accuracy. The right of the figure shows that the fine-tune accuracy of the model are not severely impacted with more harmful data mixed in the fine-tuning data. This make the fine-tuning attack even more stealthy as the fine-tuned model achieve almost the same performance for the user task.
In summary, this finding justifies that fine-tuning on data with higher ratio of harmful data can increase the harmful score of the model, resulting in the degradation of safety alignment. However, the fine-tune accuracy will not be significantly influenced even though a high ratio of harmful data is mixed in the user data.
Next we want to track two statistics: Alignment loss and hidden embedding drift to further explain the safety degradation phenomenon.
Figure 4: Alignment loss/embedding drift of SFT/non-aligned model finetuned on SST2 mixed with different ratio of harmful data. SFT model means that the pre-trained Llama2-7B is first safety aligned with SFT, while Non-Aligned means that the model is not safety aligned before user fine-tuning. Harmful ratio refers to the ratio of harmful data mixed with the fine-tuning data.
  • Alignment loss. Here we track the model’s loss on the alignment dataset (the one used for alignment) after the model is fine-tuned on fine-tuning dataset. As shown on the left of Figure 4, for the model trained using supervised fine-tuning (SFT), the alignment loss increases as the harmful data ratio rises. This suggests that the model becomes less aligned with the alignment dataset after fine-tuning with more harmful user data, implying it begins to forget its alignment knowledge. For the non-aligned model, the alignment loss starts at a high value and remains stable, even when fine-tuned with more harmful data.
  • Hidden embedding drift. To further investigate the change in alignment loss, we examine the drift in hidden embeddings after user fine-tuning, as shown on the right of Figure 4. Specifically, embedding drift is quantified as the Euclidean distance between the hidden embeddings of the aligned model (or the pre-trained model for the non-aligned case) and the fine-tuned model, all using the same alignment dataset. Here, hidden embedding refers to the output from each attention layer in the LLM. We observe that the embedding drift for the SFT model increases significantly with a higher harmful data ratio. A similar trend is seen in the non-aligned model, though the drift is less pronounced.
Figure 5: T-SNE visualization of hidden embedding drift under different harmful ratios $p$. Each point represents the embedding of each alignment data.
In summary, the above finding seems to show that hidden embedding drift over the alignment data could be a plausible reason leading to the increase of harmful score, as they share the same growing trend with the harmful score. Inspired by this finding, an alignment stage defense Vaccine [67] is proposed to mitigate this drift. We next show in Figure 5 the visualization results of the SFT and Vaccine method. As shown, When the harmful ratio is high, it is intuitive to see that the embedding of both SFT’s model and Vaccine’s model is drifting toward a specific direction. However, the embedding drift of Vaccine is slighter, making the drifted embedding still able to preserve the alignment knowledge.
With the above results, we show that how the loss over alignment data is changed with different ratio of harmful data used in harmful fine-tuning. Next we want to study how the three statistics evolved with the fine-tuning steps in the fine-tuning stage. The three statistics involve: i) Harmful score. ii) Harmful training loss, which measures the loss of those harmful data used in the fine-tuning stage, and iii) Harmful testing loss, which measures the loss of those harmful data that the model never sees in the fine-tuning stage. Here we consistently adopt harmful ratio $p=0.1$, and the same Llama2-7B model and SST2 dataset is used for evaluation.
Figure 6: Model statistics (Left: harmful score, Middle: harmful training loss, Right: harmful testing loss) after fine-tuning on pure SST2/harmful data for different steps. Specially, harmful score measures how harmful the model is (the smaller the better), harmful training loss refers to the loss over the harmful data used in fine-tuning, while harmful testing loss refers to that over the testing harmful data that the model never sees in fine-tuning stage.
Harmful score. The left of Figure 6 shows that the model’s harmful score will substantially increase along with optimization steps invested in fine-tuning on a pure harmful dataset. On the contrary, the harmful score will not be affected much via fine-tuning on a pure SST2 dataset. This demonstrates that harmful fine-tuning on harmful data is indeed more dangerous than fine-tuning on benign data.
Training/testing Harmful loss. As shown in the middle of Figure 6, more fine-tuning steps on harmful data significantly reduces the harmful training loss, which means the model starts to fit the harmful data. On the contrary, training on SST2 would only slightly increase the harmful training loss. The right of Figure 6 shows a similar trend, indicating that the fitting to the training harmful data can indeed be generalized to other unseen harmful data. This phenomenon indicates the reduction of harmful loss because of fine-tuning on harmful loss is also a plausible reason leading to safety degradation.
Alignment loss vs. Harmful loss. The previous results show that harmful fine-tuning can increase the alignment loss, leading to "forgetting" of alignment knowledge. At the same time, we show that harmful fine-tuning can decrease the harmful loss, leading to "revitalization" of harmful knowledge. Based on this finding, there are two directions for mitigation strategy design. The first direction is how to better preserve the alignment knowledge, such that it is harder to be disturbed, as represented by Vaccine [67]. Lisa [66] and T-Vaccine [96]. Under the assumption that we have in-distributional harmful data for defense design, another direction of defense aim to remove the harmful knowledge (e.g., RepNoise[127], Antidote [63]) or attenuate harmful knowledge’s impact on the model (e.g., TAR [143], Booster [65]). Efforts should be make in both the two directions to mitigate the impact of harmful fine-tuning to achieve either one of the two ultimate goals: i) harmful knowledge can be sufficiently learned, (i.e., the harmful training loss is able to reduced to 0). However, the alignment is well-preserved such that the harmful knowledge cannot be generalized to other unseen harmful question (i.e., alignment loss is still 0, which means after learning harmful knowledge, model is still able to give refusal answer to other harmful questions.). ii) The harmful training loss is very high even after fine-tuning (i.e., those harmful data are untrainable).
@@ 4 Existing Attacks and Defenses
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [LLM Harmful Fine-tuning , fill=gray!45, parent [Harmful fine-tuning Attack, perception [Vanilla, perception [Shadow Alignment[166], Yi et al. [169], Lermen et al. [84], Zhan et al. [176], Chen et al. [20], Hawkins et al. [49], Poppi et al. [119] , perception_work ] ] [Advanced, perception [Bi-directional Anchoring [51], Covert Malicious Fine-tuning [48], perception_work ] ] ] [Harmful Fine-tuning Defense, perception [Alignment-stage, perception [Optimization, perception [Vaccine [67], RepNoise[127], TAR [143], Booster [65], RSN-Tune [1], T-Vaccine [96] , perception_work] ] [Data Augmentation, perception [CTRL [100], perception_work] ] ] [Fine-tuning-stage, perception [Distance Regularization, perception [LDIFS [108], Freeze [159], Constrain-SFT [121], Paraphrase [39], ML-LR [38], Freeze+ [4], SaLoRA [5], , perception_work] ] [Data Augmentation, perception [SafeInstr [11], VLGuard [182], Lisa [66] , perception_work] ] [Prompt Engineer, perception [BEA[156], PTST [102], , perception_work] ] [Detection/Filtration, perception [Seal [132], SAFT [23] , perception_work] ] ] [Post-fine-tuning stage, perception [LAT [17], SOMF [171], Safe LoRA [53], Antidote [63], SafetyLock [181], , perception_work] ] ] [Mechanism Analysis, perception [On Attacks, perception [Leong et al. [83], Peng et al. [116], Anonymous [7], Guo et al. [46], perception_work] ] [On Defenses, perception [ Anonymous [3] , perception_work] ] ] ]
Figure 7: Mind map for the existing literature on harmful fine-tuning attacks.
In this section, we summarize the current research progress on harmful fine-tuning attacks and defenses. A mindmap is shown in Figure 7 for summarization.
@@ 4.1 Harmful Fine-tuning Attacks
Shadow Alignment[166] 100 malicious examples can subvert alignment Shawdow alignment SFT (full) Oct 4, 2023
@@
Hawkins et al. [49] fine-tuning on developer-tuned models with benign data compromises safety Dolly SFT (LoRA) Oct 21, 2024